# Rice-like complexity lower bounds for Boolean and uniform automata networks

### v1

@@ Abstract
Automata networks are a versatile model of finite discrete dynamical systems composed of interacting entities (the automata), able to embed any directed graph as a dynamics on its space of configurations (the set of vertices, representing all the assignments of a state to each entity). In this world, virtually any question is decidable by a simple exhaustive search. We lever the Rice-like complexity lower bound, stating that any non-trivial monadic second order logic question on the graph of its dynamics is ${\mathsf{NP}}$-hard or ${\mathsf{coNP}}$-hard (given the automata network description), to bounded alphabets (including the Boolean case). This restriction is particularly meaningful for applications to “complex systems”, where each entity has a restricted set of possible states (its alphabet). For the non-deterministic case, trivial questions are solvable in constant time, hence there is a sharp gap in complexity for the algorithmic solving of concrete problems on them. For the non-deterministic case, non-triviality is defined at bounded treewidth, which offers a structure to establish metatheorems of complexity lower bounds.
@@ 1 Introduction
A pillar of computer science is Rice theorem [30]: any non-trivial semantic property of programs is undecidable. It is striking for its generality and the sharp dichotomy of difficulty between trivial and non-trivial problems. We look for analogs in this spirit. A finite discrete dynamical system can only ask decidable questions (simply because it is finite), hence we shift the perspective from computability to complexity theory. We lever the Rice-like complexity lower bounds of [9, 10], from succinct graph representations (where it contrasts Courcelle theorem) to natural models of interacting entities (automata networks on bounded alphabets). Such metatheorems are obtained using the expressiveness of graph logics, all at monadic second order (MSO) in the present work. A key consists in defining a notion of non-triviality as general as possible. It turns out to be a sharp and deep dichotomy in the deterministic case (AN) because trivial questions are answered in constant time. In the non-deterministic case (NAN), we employ a notion of non-triviality relative to dynamics of bounded treewidth, which we call arborescence (its necessity is addressed in [10], and discussed in the perspectives). Our results apply to uniform automata networks, where all automata are constrained to have the same alphabet size (denoted $q$). All the necessary concepts will be defined in the preliminary section. We gather our main results in one statement, where in bold is the problem consisting in deciding whether the dynamics of an automata network (given a description of the behavior of its entities) verifies some property $\psi$ expressed in graph logics.
@@
c. For any arborescent MSO formula, $\psi$-NAN-dynamics is ${\mathsf{NP}}$- or ${\mathsf{coNP}}$-hard.
d. For any $q$-arborescent MSO formula, $\psi$-$q$-NAN-dynamics is ${\mathsf{NP}}$- or ${\mathsf{coNP}}$-hard.
Part c of Theorem 1 is proven in [10]. The ${\mathsf{NP}}$ and ${\mathsf{coNP}}$ symmetry is necessary in such a general statement, because some problems expressible at first order (such as the existence of a fixed point in the dynamics) are known to be ${\mathsf{NP}}$-complete [1], and the graph logics we consider are closed by complemention (simply by adding a negation on top of the formula). The proof technique nevertheless gives a clear characterization of which of these two lower bounds is proven for each formula $\psi$, through the concept of saturating graph (denoted $\Omega_{m}$ because it only depends on the quantifier rank $m$ of $\psi$) introduced in [10]. If $\Omega_{m}$ satisfies an MSO formula $\psi$, then the corresponding $\psi$-dynamics problem is proven to be ${\mathsf{NP}}$-hard, otherwise it is proven to be ${\mathsf{coNP}}$-hard.
In Section 2 we introduce all the necessary notions involved in the statement and in the proof of Theorem 1, including graph logics and tree-decompositions. Theorem 1 extends known results in two directions:
@@
The proof technique developed in [10], on which our extensions are grounded, is detailed in Section 3 where all the ingredients to design a polytime reduction from SAT are exposed. In Section 4 we apply the technique to deterministic dynamics, i.e. to graphs of out-degree $1$, and prove part a of Theorem 1. Section 5 is dedicated to the study of arithmetical considerations in order to obtain $q$-uniform dynamics (on a configuration space of size $q^{n}$ for some number $n$ of automata) when pumping a part of size $a$ from a graph of size $b$ (that is, obtaining potential dynamics of size $ak+b$ where $k$ is the number of copies of size $a$ which have been added to the initial graph of size $b$). We give a separate consideration to the Boolean case $q=2$, which is simpler and gives the intuitions for the last ingredients of our main results. In Section 6 we apply the arithmetics of the previous section to derive parts b and d of Theorem 1 which, to our point of view, are the most impacting new results presented in this paper, already asked for at the end of [9]. We conclude and present perspectives on how to pursue the quest for metatheorems on the complexity of discrete dynamical systems in Section 7.
@@ 2 Preliminaries
For $n,q\in\mathbb{N}_{+}$, we denote $[n]=\{1,\dots,n\}$ and $\llbracket q\rrbracket=\{0,\dots,q-1\}$. Given a directed graph (digraph) $G=(V(G),E(G))$, we consider its size as $|G|=|V(G)|$. Given a vertex $v\in V(G)$, the out-neighbors of $v$ are denoted $G(v)=\{u\mid(v,u)\in E(G)\}$, and $|G(v)|$ is the out-degree of $v$. We say that a graph $G$ is of out-degree (exactly) $d\in\mathbb{N}$ when all its vertices have out-degree $d$.
@@ 2.1 Automata networks
A non-deterministic automata network (NAN) of size $n$ is a relation $f\subseteq X\times X$ where $f(x)$ is the set of possible transitions from configuration $x\in X$ (again with $X=\prod_{i\in[n]}A_{i}$). Remark that defining local relations $(f_{i}\subseteq X\times A_{i})_{i\in[n]}$ is not always possible, because it would suggest that $f(x)=\{y\mid\forall i\in[n],y_{i}\in f_{i}(x)\}$ but this is not possible for all relations $f$ when $n>1$ (and the number of automata is at the heart of our considerations, in particular regarding $q$-uniformity). The non-determinism we consider in this paper is global, by contrast with local non-determinism defined through local relations, where the image of a configuration is taken among the Cartesian product of the possibilities for each automaton ($f(x)\in f_{1}(x)\times f_{2}(x)\times\dots\times f_{n}(x)$).
When all the state sets $A_{i}$ have the same size $q$, we say that $f$ is a $q$-uniform (non-deterministic) automata network ($q$-AN or $q$-NAN), For $q=2$ we have $X=\{0,1\}^{n}$ and we call $f$ a Boolean (non-deterministic) automata network (BAN or BNAN).
@@
Figure 2: Example non-deterministic $q$-uniform automata network of size $n=2$ on alphabet $\llbracket q\rrbracket=\{0,1,2\}$ with $q=3$. Circuit of the relation $f\subseteq\llbracket q\rrbracket^{n}\times\llbracket q\rrbracket^{n}$ (left) and transition digraph dynamics $\mathcal{G}_{f}$ on configuration space $\llbracket q\rrbracket^{n}$ (right). A configuration is encoded on four bits, where inputs $0000$ to $0001$ correspond respectively to configurations $00$ to $22$. The doubly circled node represents the evaluation of propositionnal formula $S=(x_{1}\vee x_{2})\wedge(\neg x_{1}\vee\neg x_{2})\wedge x_{3}$ on the three first bits of the circuit’s input.
@@
Given our focus on computational complexity, it is important to explicit how ANs and NANs are encoded, to be given as inputs to problems. An AN $f$ is encoded as one Boolean circuit with $\lceil\log_{2}|X|\rceil$ input bits and $\lceil\log_{2}|X|\rceil$ output bits. The state set sizes $(|A_{i}|)_{i\in[n]}$ are also part of the input, to encode/decode configurations as bitstrings. Given a configuration $x\in X$ as input to the circuit, its image $f(x)$ is read as the output taken modulo $|X|$ (to avoid the ${\mathsf{coNP}}$-complete problem of checking that the circuit always outputs a bitstring from 00 to $|X|-1$). In the Boolean case, each input/output bit of the circuit is the state of an automaton (circuits generalize propositional formulas, but all our results still hold for an encoding through formulas). A NAN $f$ is encoded as one Boolean circuit with $2\cdot\lceil\log_{2}|X|\rceil$ input bits and $1$ output bits, indicating whether the two input configurations $x$ and $y$ verify $y\in f(x)$. The circuit size of an AN is restricted to some $\mathcal{O}(|X|\log_{2}|X|)$ (the size of a lookup table for $\mathcal{G}_{f}$ of out-degree $1$), and that of a NAN to some $\mathcal{O}(|X|^{2})$ (the size of an adjacency matrix for $\mathcal{G}_{f}$). These encodings are also called succinct graph representations of $\mathcal{G}_{f}$. When constructing ANs or NANs in our proofs, it will always be straightforward to convert (in polynomial time) our descriptions of dynamics into local functions and their encodings as circuits. Examples of AN and NAN are given on Figures 1 and 2, respectively.
@@ 2.2 Graph Logics
[…]
We will express general complexity lower bounds for non-trivial fromulas, but this notion depends on the ANs or NANs under consideration. A formula $\psi$ is non-trivial when it has infinitely many models and infinitely many counter-models, among graphs of out-degree $1$. A formula $\psi$ is arborescent when there exists $k$ such that $\psi$ has infinitely many models of treewidth at most $k$ and infinitely many counter-models of treewidth at most $k$ (the definition of treewidth is recalled below). According to Remark 1, these notions respectively correspond to deterministic ($\mathcal{G}_{f}$ for ANs) and non-deterministic ($\mathcal{G}_{f}$ for NANs) dynamics. When considering only graphs $\mathcal{G}_{f}$ for $q$-uniform networks, we obtain the notions of $q$-non-trivial and $q$-arborescent formulas.
@@
Input: circuit of an AN $f$ of size $n$.
@@
Input: circuit of an NAN $f$ of size $n$.
@@
  • $\exists x,x\to x\wedge(\forall y,y\neq x\implies(\exists z,y\neq z\wedge y\to z\wedge z\to y))$ requires $|V(\mathcal{G}_{f})|$ to be odd, hence it is non-trivial in general, but trivial in the Boolean case.
@@
  • $\forall x,\exists y,x\to y\wedge(\forall z,z\neq y\implies\neg(x\to z))$ means that $\mathcal{G}_{f}$ has out-degree $1$, which is trivial on determistic dynamics.
Most formulas above do not use the second order quantifier, hence they are First Order (FO) formulas. It is known that properties such as connecitivity, acyclicity, 2-colorability and planarity are expressible in MSO, but not in FO [19, Chapter 3].
@@
2.3 Boundaried graphs and equivalence classes
Let $k\in\mathbb{N}$. A $k$-boundaried graph is a digraph $G$ equipped with a sequence $P(G)=(p_{1},...,p_{k})$ of $k$ distinct vertices from $V(G)$, called ports. Given two $k$-boundaried graphs $G$, $G^{\prime}$ with ports $P(G)=(p_{1},...,p_{k})$ and $P(G^{\prime})=(p^{\prime}_{1},...,p^{\prime}_{k})$, we define their gluing as $G\oplus G^{\prime}=H$ with:
$V(H)$ $=V(G)\cup(V(G^{\prime})\backslash P(G^{\prime})),$
$E(H)$ $=E(G)\cup E(G^{\prime}[P(G)\leftarrow P(G^{\prime})]),$
$P(H)$ $=P(G),$
where $G^{\prime}[P(G)\leftarrow P(G^{\prime})]$ denotes the graph $G^{\prime}$ where port $p^{\prime}_{i}$ is renamed $p_{i}$ (for $i\in[k]$). In words, we take the union of $G$ and $G^{\prime}$, then merge port $p_{i}$ of $G$ with port $p^{\prime}_{i}$ of $G^{\prime}$, while keeping all the edges. Observe that $\oplus$ is, strickly speaking, not commutative in general (because we keep the port vertices from the left graph, although $G\oplus G^{\prime}$ is isomorphic to $G^{\prime}\sqcup G$), but for $k=0$ we have $G\oplus G^{\prime}=G\sqcup G^{\prime}$.
A $k$-biboundaried graph is a digraph $G$ equipped with two sequences of $k$ ports: the primary ports $P_{1}(G)$, and the secondary ports $P_{2}(G)$. We still have $P_{i}(G)\subseteq V(G)$ for $i\in\{1,2\}$, but we do not necessarily have $P_{1}(G)\cap P_{2}(G)=\emptyset$. Given two $k$-biboundaried graphs $G$, $G^{\prime}$, considering them as $k$-boundaried with $P(G)=P_{2}(G)$ and $P(G^{\prime})=P_{1}(G^{\prime})$, and denoting $H=G\oplus G^{\prime}$ (for the $k$-boundaried graphs), we define the gluing of the two $k$-biboundaried graphs as $G\oplus G^{\prime}=H^{\prime}$ with
$V(H^{\prime})$ $=V(H),$
$E(H^{\prime})$ $=E(H),$
$P_{1}(H^{\prime})$ $=P_{1}(G),$
$P_{2}(H^{\prime})$ $=P_{2}(G^{\prime}).$
In words, we take the union of $G$ and $G^{\prime}$, then merge ports $P_{2}(G)$ with $P_{1}(G^{\prime})$. The operation $\oplus$ is not commutative in general, although it still corresponds to the disjoint union for $k=0$.
As the reader can observe, we overload $\oplus$ and it becomes unclear whether it applies to boundaried or biboundaried graphs. To prevent ambiguities, we consider a boundaried graph $G$ as biboundaried with $P_{1}(G)=P_{2}(G)=P(G)$ (this conversion commutes with $\oplus$). A $k$-biboundaried graph is called a $k$-graph, for short. The operator $\oplus$ only applies to graphs with the same number $k$ of ports. Primary (resp. secondary) ports are not necessary to obtain a digraph from right (resp. left) gluing, therefore we will sometimes omit them. See Figure 3 for an example.
Figure 3: Example biboundaried graphs and their gluing. Primary ports are represented by hatched nodes, and secondary ports by grey nodes. A graph $G$ with $P_{1}(G)=(1,3)$ and $P_{2}(G)=(4,5)$ (left). A graph $G^{\prime}$ with $P_{1}(G^{\prime})=(7,6)$ and $P_{2}(G^{\prime})=(10,11)$ (middle). The graph $G\oplus G^{\prime}=H$ with $P_{1}(H)=(1,3)$ and $P_{2}(H)=(10,11)$ (right).
Now let $\Gamma=\{G_{i}\}_{i\in\mathcal{I}}$ be a finite family of $k$-graphs. For $\omega=\omega_{1}\dots\omega_{\mu}$ a nonempty word over alphabet $\mathcal{I}$, the operation $\Delta^{\Gamma}(\omega)$ is defined by induction on the length $\mu\in\mathbb{N}_{+}$ of $\omega$ as follows:
$\Delta^{\Gamma}(\omega_{1})=G_{\omega_{1}},\qquad\Delta^{\Gamma}(\omega_{1}\dots\omega_{\mu})=\Delta^{\Gamma}(\omega_{1}\dots\omega_{\mu-1})\oplus G_{\omega_{\mu}}.$
Let $G,G^{\prime}$ be $k$-graphs. Given an MSO sentence $\psi$, we write $G\sim_{k,\psi}G^{\prime}$ if and only if for every $k$-graph $H$, we have: $G\oplus H\vDash\psi\iff G^{\prime}\oplus H\vDash\psi$.
Denote $\Sigma_{k,\psi}$ the set of equivalence classes of $\sim_{k,\psi}$ for $k$-graphs. By [5, Theorem 13.1.1], if $k$ and $\psi$ are fixed then $\Sigma_{k,\psi}$ is finite.
We may drop subscripts $k$ and/or $\psi$ when they are clear from the context.
2.4 Tree-decompositions and treewidth
All our graphs will be directed. The subtree of a tree $T$ rooted in vertex $v\in V(T)$ is the induced subgraph obtained from the descendants of $v$ (the root of the subtree).
A tree-decomposition of a digraph $G=(V,E)$ is a tree $T$ whose nodes are labelled with ordered (to become ports) subsets of $V$, called bags, satisfying the three conditions below. If $v$ is a node of $T$, we write $B_{T}(v)\subseteq V$ the corresponding bag (when $T$ is clear from the context, we drop the subscript notation).
  1. 1.
    Every vertex $x$ of $G$ is containted in at least one bag $B(v)$ of $T$.
  2. 2.
    For every edge $(x,y)$ of $G$, both vertices $x$ and $y$ are contained in at least one bag $B(v)$ of $T$.

### v2

@@ Abstract
Automata networks are a versatile model of finite discrete dynamical systems composed of interacting entities (the automata), able to embed any directed graph as a dynamics on its space of configurations (the set of vertices, representing all the assignments of a state to each entity). In this world, virtually any question is decidable by a simple exhaustive search. We lever the Rice-like complexity lower bound, stating that any non-trivial monadic second order logic question on the graph of its dynamics is $\mathsf{NP}$}-hard or $\mathsf{coNP}$}-hard (given the automata network description), to bounded alphabets (including the Boolean case). This restriction is particularly meaningful for applications to “complex systems”, where each entity has a restricted set of possible states (its alphabet). For the deterministic case, trivial questions are solvable in constant time, hence there is a sharp gap in complexity for the algorithmic solving of concrete problems on them. For the non-deterministic case, non-triviality is defined at bounded cliquewidth, which offers a structure to establish metatheorems of complexity lower bounds.
@@ 1 Introduction
A pillar of computer science is Rice theorem [32]: any non-trivial semantic property of programs is undecidable. It is striking for its generality and the sharp dichotomy of difficulty between trivial and non-trivial problems. We look for analogs in this spirit. A finite discrete dynamical system can only raise decidable questions (simply because it is finite), hence we shift the perspective from computability to complexity theory. We lever the Rice-like complexity lower bounds of [11, 10], from succinct graph representations (where it contrasts Courcelle theorem) to natural models of interacting entities (automata networks on bounded alphabets). Such metatheorems are obtained using the expressiveness of graph logics, all at monadic second order (MSO) in the present work. A key consists in defining a notion of non-triviality as general as possible. It turns out to be a sharp and deep dichotomy in the deterministic case (AN) because trivial questions are answered in constant time. In the non-deterministic case (NAN), we employ a notion of non-triviality relative to dynamics of bounded cliquewidth, which we call cw-non-trivial (its necessity is addressed in [10] and discussed in the perspectives). Our results apply to uniform automata networks where all automata are constrained to have the same alphabet size (denoted $q$). All the necessary concepts will be defined in the preliminary section. We gather our main results in one statement, where in bold is the problem consisting in deciding whether the dynamics of an automata network (given a description of the behavior of its entities) verifies some property $\psi$ expressed in graph logics.
@@
c. For any cw-non-trivial MSO formula, $\psi$-NAN-dynamics is $\mathsf{NP}$}- or $\mathsf{coNP}$}-hard.
d. For any $q$-cw-non-trivial MSO formula, $\psi$-$q$-NAN-dynamics is $\mathsf{NP}$}- or $\mathsf{coNP}$}-hard.
Part c of Theorem 1 is proven in [10]. The $\mathsf{NP}$} and $\mathsf{coNP}$} symmetry is necessary in such a general statement, because some problems expressible at first order (such as the existence of a fixed point in the dynamics) are known to be $\mathsf{NP}$}-complete [1], and the graph logics we consider are closed by complementation (simply by adding a negation on top of the formula). The proof technique nevertheless gives a clear characterization of which of these two lower bounds is proven for each formula $\psi$, through the concept of saturating graph (denoted $\Omega_{m}$ because it only depends on the quantifier rank $m$ of $\psi$) introduced in [10]. If $\Omega_{m}$ satisfies an MSO formula $\psi$, then the corresponding $\psi$-dynamics problem is proven to be $\mathsf{NP}$}-hard, otherwise it is proven to be $\mathsf{coNP}$}-hard.
In Section 2 we introduce all the necessary notions involved in the statement and in the proof of Theorem 1, including graph logics and clique-decompositions. Theorem 1 extends known results in two directions:
@@
The proof technique developed in [10], on which our extensions are grounded, is detailed in Section 3 where all the ingredients to design a polytime reduction from SAT are exposed. In Section 4 we apply the technique to deterministic dynamics, i.e. to graphs of out-degree $1$, and prove Part a of Theorem 1. Section 5 is dedicated to the study of arithmetical considerations in order to obtain $q$-uniform dynamics (on a configuration space of size $q^{n}$ for some number $n$ of automata) when pumping a part of size $a$ from a graph of size $b$ (that is, obtaining potential dynamics of size $ak+b$ where $k$ is the number of copies of size $a$ which have been added to the initial graph of size $b$). We give a separate consideration to the Boolean case $q=2$, which is simpler and gives the intuitions for the last ingredients of our main results. In Section 6 we apply the arithmetics of the previous section to derive Parts b and d of Theorem 1 which, to our point of view, are the most impacting new results presented in this paper, already asked for at the end of [11]. We conclude and present perspectives on how to pursue the quest for metatheorems on the complexity of discrete dynamical systems in Section 7.111This work is an update of [15] (from v1 to v2) in order to bring the results from dynamics of bounded treewidth to dynamics of bounded cliquewidth, according to a recent update of [10] performing this extension of the results (from v2 to v3).
@@ 2 Preliminaries
For $n,q\in\mathbb{N}_{+}$, we denote $n]=\{1,\dots,n\}$ and $\llbracket q\rrbracket=\{0,\dots,q-1\}$. Given a directed graph (digraph) $G=(V(G),E(G))$, we consider its size as $|G|=|V(G)|$. Given a vertex $v\in V(G)$, the out-neighbors of $v$ are denoted $G(v)=\{u\mid(v,u)\in E(G)\}$, and $|G(v)|$ is the out-degree of $v$. We say that a graph $G$ is of out-degree (exactly) $d\in\mathbb{N}$ when all its vertices have out-degree $d$.
@@ 2.1 Automata networks
A non-deterministic automata network (NAN) of size $n$ is a relation $f\subseteq X\times X$ where $f(x)$ is the set of possible transitions from configuration $x\in X$ (again with $X=\prod_{i\in[n]}A_{i}$). Remark that defining local relations $(f_{i}\subseteq X\times A_{i})_{i\in[n]}$ is not always possible, because it would suggest that $f(x)=\{y\mid\forall i\in[n],y_{i}\in f_{i}(x)\}$ but this is not possible for all relations $f$ when $n>1$ (and the number of automata is at the heart of our considerations, in particular regarding $q$-uniformity). The non-determinism we consider in this paper is global, by contrast with local non-determinism defined through local relations, where the image of a configuration $x$ is taken among the Cartesian product of the possibilities for each automaton ($f(x)\in f_{1}(x)\times f_{2}(x)\times\dots\times f_{n}(x)$).
When all the state sets $A_{i}$ have the same size $q$, we say that $f$ is a $q$-uniform (non-deterministic) automata network ($q$-AN or $q$-NAN). For $q=2$ we have $X=\{0,1\}^{n}$ and we call $f$ a Boolean (non-deterministic) automata network (BAN or BNAN).
@@
Figure 2: Example non-deterministic $q$-uniform automata network of size $n=2$ on alphabet $\llbracket q\rrbracket=\{0,1,2\}$ with $q=3$. Circuit of the relation $f\subseteq\llbracket q\rrbracket^{n}\times\llbracket q\rrbracket^{n}$ (left) and transition digraph dynamics $\mathcal{G}_{f}$ on configuration space $\llbracket q\rrbracket^{n}$ (right). A configuration is encoded on four bits, where inputs $0000$ to $0001$ correspond respectively to configurations $00$ to $22$. The doubly circled node represents the evaluation of propositional formula $S=(x_{1}\vee x_{2})\wedge(\neg x_{1}\vee\neg x_{2})\wedge x_{3}$ on the three first bits of the circuit’s input.
@@
Given our focus on computational complexity, it is important to make explicit how ANs and NANs are encoded when given as inputs to problems. An AN $f$ is encoded as one Boolean circuit with $\lceil\log_{2}|X|\rceil$ input bits and $\lceil\log_{2}|X|\rceil$ output bits. The state set sizes $(|A_{i}|)_{i\in[n]}$ are also part of the input in order to encode/decode configurations as bitstrings, but formally we only require to provide the global size $N=|X|$. Given a configuration $x\in X$ as input to the circuit, its image $f(x)$ is read as the output taken modulo $N$ (to avoid the $\mathsf{coNP}$}-complete problem of checking that the circuit always outputs a bitstring from 0 to N1N-1). In the Boolean case, each input/output bit of the circuit is the state of an automaton (circuits generalize propositional formulas, but all our results still hold for an encoding through formulas). A NAN $f$ is encoded as one Boolean circuit with $2\cdot\lceil\log_{2}|X|\rceil$ input bits and $1$ output bit, indicating whether the two input configurations $x$ and $y$ verify $y\in f(x)$. The circuit size of an AN is restricted to some $\mathcal{O}(|X|\log_{2}|X|)$ (the size of a lookup table for $\mathcal{G}_{f}$ of out-degree $1$), and that of a NAN to some $\mathcal{O}(|X|^{2})$ (the size of an adjacency matrix for $\mathcal{G}_{f}$). These encodings are also called succinct graph representations of $\mathcal{G}_{f}$ [9]. It is important to precise the size $N=|X|$ of the configuration space when there is not restriction on the alphabets, but for $q$-uniform networks it is enough to give the number $n$ of automata (then we deduce that there are $q^{n}$ configurations). When constructing ANs or NANs in our proofs, it will always be straightforward to convert (in polynomial time) our descriptions of dynamics into local functions and their encodings as circuits. Examples of AN and NAN are given on Figures 1 and 2, respectively.
@@ 2.2 Graph logics
If $P$ is a property that automata networks may or may not satisfy, and $f$ is an AN or a NAN, then we write $f\vDash P$ if $f$ satisfies $P$, and $f\nvDash P$ otherwise. We say that $f$ is a model of $P$ in the first case, and a counter-model otherwise. This is an abuse of notation, and we need to know the exact nature of $P$ to know its precise meaning. In particular, we will study properties expressible on graph monadic second order logic (MSO) over the signature $\{=,\to\}$, where $\to$ is a binary relation such that $x\to y$ if and only if $y=f(x)$ or $y\in f(x)$ (that is, $(x,y)\in E(\mathcal{G}_{f})$). Our formulas express graphical properties of $\mathcal{G}_{f}$, with logical operators ($\wedge$, $\vee$, $\neg$, $\implies$), and quantifications ($\exists,\forall$) on vertices (configurations) or sets of vertices (with the addition of $\in$ to the signature, for set membership). The quantifier rank of a formula $\psi$ is its depth of quantifier nesting. If $G$ and $G^{\prime}$ are two structures, we write $G\equiv_{m}G^{\prime}$ when they satisfy the same formulas of quantifier rank $m$. We write $\psi\equiv\psi^{\prime}$ when the two formulas $\psi$ and $\psi^{\prime}$ have the same models and the same counter-models.
We will express general complexity lower bounds for non-trivial formulas, a notion which depends on the family of ANs or NANs under consideration. A formula $\psi$ is non-trivial when it has infinitely many models and infinitely many counter-models, among graphs of out-degree $1$. A formula $\psi$ is cw-non-trivial when there exists $k$ such that $\psi$ has infinitely many models of cliquewidth at most $k$ and infinitely many counter-models of cliquewidth at most $k$ (the definition of cliquewidth is recalled below). According to Remark 1, these notions respectively correspond to deterministic ($\mathcal{G}_{f}$ for ANs) and non-deterministic ($\mathcal{G}_{f}$ for NANs) dynamics. When considering only graphs $\mathcal{G}_{f}$ for $q$-uniform networks, we obtain the notions of $q$-non-trivial and $q$-cw-non-trivial formulas.
@@
Input: circuit of an AN $f$ with $N$ configurations.
@@
Input: circuit of an NAN $f$ with $N$ configurations.
@@
  • $\exists x,x\to x\wedge(\forall y,y\neq x\implies(\exists z,y\neq z\wedge y\to z\wedge z\to y))$ requires $|V(\mathcal{G}_{f})|$ to be odd, hence it is non-trivial in general, but trivial in the Boolean case ($2$-trivial).
@@
  • $\forall x,\exists y,x\to y\wedge(\forall z,z\neq y\implies\neg(x\to z))$ means that $\mathcal{G}_{f}$ has out-degree $1$, which is trivial on deterministic dynamics (but it is cw-non-trivial).
Most formulas above do not use the second order quantifier, hence they are first order (FO) formulas. It is known that properties such as connectivity, acyclicity, 2-colorability and planarity are expressible in MSO, but not in FO [21, Chapter 3].
Observe that the signature $\{=,\to\}$ only allows to express properties up to isomorphism, that is if $\mathcal{G}_{f}$ and $\mathcal{G}_{g}$ are isomorphic then $\mathcal{G}_{f}\equiv_{m}\mathcal{G}_{g}$ for all $m$.
2.3 Clique-decompositions and cliquewidth
For a finite set of $k$ colors $C=\{1,\dots,k\}$, a $k$-colored graph $G=(V,E,C_{G})$ is a (directed) graph where each vertex is assigned one color by $C_{G}:V\to C$. The four following operations are used to construct $k$-colored graphs:
  1. 1.
    $\mathrm{constant}_{i}$ (arity 0): given a color $i\in C$, it is the graph $G=(\{v\},\emptyset,C_{G})$ with $C_{G}(v)=i$ for some fresh vertex label $v$;
  2. 2.